OpenStudy (anonymous):

Please Help Me W/ 2 Questions ... Medal ((&)) Fan Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). A) y = x + 8 B) y = x + 3 C) y = -x + 8 D) y = -x + 3 ************ Select the coordinates of two points on the line y = -2. A) (2, -2) and (-2, 2) B) (-2, -2) and (-2, 0) C) (2, -2) and (0, -2) D) (-2, 2) and (-2, -2) **************

OpenStudy (anonymous):

Number 2 is C and let me look at number 1

OpenStudy (anonymous):

Ok but are you sure about C

OpenStudy (anonymous):

@charlotte123

OpenStudy (anonymous):

Yes, because it says it's on the y line and that it's on -2, the y is the second number in a set so both would have to have -2 as the second number

OpenStudy (anonymous):

ok ((&)) answer choices for #1 are.... y = 4/3x + 8 y = 4/3x + 3 y = -3/4x + 8 y = -3/4x + 3

Elsa213 (elsa213):

Write an equation for the line parallel to the line `4x+3y=9` and passing through the point (-6,2) Parallel lines have the same slope. Writing the given line in slope-intercept form we have:...

OpenStudy (anonymous):

A

Elsa213 (elsa213):

NO ANSWERS MISTER!

OpenStudy (charlotte123):

Thanks a bunch cecor

Elsa213 (elsa213):

:O @charlotte123 u allow dis?!?!?!

OpenStudy (anonymous):

thx @cecormier but i was trying to learn how to do it

OpenStudy (charlotte123):

*bans*

OpenStudy (anonymous):

calm down you all

OpenStudy (charlotte123):

"calm down you all"

OpenStudy (charlotte123):

Ma'm I'm gonna have to ask you to calm down before I ban you too pls

OpenStudy (anonymous):

umm ban me for what

OpenStudy (charlotte123):

YOU HAVE THE RIGHT TO REMAIN OWL MA'M

OpenStudy (anonymous):

Girl bye idgaf if you ban me or not you aint god i can care less about your ban im still living and breathing i dont need you df

OpenStudy (charlotte123):

I'm dying

OpenStudy (charlotte123):

L O L

OpenStudy (anonymous):

Dismissed

Elsa213 (elsa213):

:O TAKE HER DOWN CHAR!!!!! *Grabs stick and starts hitting brala*

OpenStudy (anonymous):

this is stupid yall petty

OpenStudy (charlotte123):

L O L

OpenStudy (anonymous):

@charlotte123 You have been dismissed by @brala00 Goodbye!

OpenStudy (anonymous):

anyways thx @cecormier

OpenStudy (charlotte123):

ROOOOOOOOOOOOOOOOOOOLFOFLOFLFOLOFLOFLOFLFFLOFLFOFL

OpenStudy (anonymous):

Frfr she started w/ me acting like she queen of england @Jesus.

OpenStudy (charlotte123):

"queen of england" That's queen of the world to you

OpenStudy (charlotte123):

mhmmmmmmmmmmmm

Elsa213 (elsa213):

:O U have du right 2 silence Ma'am! U dont start with my queen >_>

OpenStudy (charlotte123):

she wasn't readyyyyyyyyyyyyyyyyyyyyyyy

Elsa213 (elsa213):

XD

OpenStudy (anonymous):

Sharkeisha No! She wasnt ready!

OpenStudy (anonymous):

girl goodbye i got work to do i didnt come on openstudy to argue with someone i know nothing about if you have a problem you can get off my question and go about your business

OpenStudy (charlotte123):

I'm dying guys

OpenStudy (charlotte123):

Brala I loe you <3

OpenStudy (anonymous):

Its love * check your grammar

OpenStudy (anonymous):

http://prntscr.com/4b7awi

OpenStudy (charlotte123):

Brala my stew Brala my shoe

Elsa213 (elsa213):

we were playing brala

OpenStudy (charlotte123):

It's loe

OpenStudy (charlotte123):

Guys http://prntscr.com/4b7awi I can't anymore

OpenStudy (anonymous):

well im not exactly nice and i dont like to be played w/ like that so i appreciate if you stop and leave me alone @Elsa213

OpenStudy (charlotte123):

Wellllllllllllllllllll obviouslyyyyyyyyyyyyyyyyyyy

Elsa213 (elsa213):

:o okay ma'am....... u got 2 chill those diapers down ma'am.

OpenStudy (anonymous):

bye prepare to get blocked

OpenStudy (charlotte123):

Oooooooooooooooo I'ma get blocked bby don't I loe you </3

OpenStudy (anonymous):

and no df i dont because i dont have time for jokes either you help me or leave me df alone go watch kevin hart if you wanna find something to laugh at

Elsa213 (elsa213):

ROOOOOOOOOOOOOOOOOOFLOFLOFLOFLOFLOFLOFL

OpenStudy (charlotte123):

ROLFOFLFOLFOLFOFLFOLFOFLOFLFOLFOFLOFLFOLFOFLF

OpenStudy (charlotte123):

HOLY ₍₍ ◝(・ω・)◟ ⁾⁾ I'm done

OpenStudy (bibby):

Why would I watch Kevin Hart if I wanted to laugh?

OpenStudy (charlotte123):

#ded

Elsa213 (elsa213):

agree

OpenStudy (tester97):

@brala00 are you jews?

OpenStudy (charlotte123):

must be

Elsa213 (elsa213):

i think so

OpenStudy (charlotte123):

But that's none of our business http://prntscr.com/4b7awi

Elsa213 (elsa213):

she/he created new question...... should we go der?

OpenStudy (charlotte123):

#yolo

OpenStudy (tester97):

@Gin_Ichimaru